If Fn is the n-th Fibonacci number, show that lim "\lim_{n \to \infty} \frac { F_{n+1}}{F_n} = \frac{1+\sqrt{5}}{2}"





1
Expert's answer
2021-11-30T16:06:23-0500

Prove that


"F_n=\dfrac{1}{\sqrt{5}}\bigg((\dfrac{1+\sqrt{5}}{2})^{n}-(\dfrac{1-\sqrt{5}}{2})^{n}\bigg)"


Let’s  look at the interesting quadratic


"x^2-x-1=0"

The roots of this quadratic are

"x=\dfrac{1\pm\sqrt{5}}{2}"

This quadratic can also be written as


"x^2=x+1"

From this, we can write expressions for all "x^n:"


"x=x"




"x^2=x+1"

"x^3=x\cdot x^2=x(x+1)""=x^2+x=x+1+x=2x+1"


"x^4=x\cdot x^3=x(2x+1)""=2x^2+x=2x+2+x=3x+2"

"x^5=5x+3=x^3+x^4"

"x^6=8x+5=x^4+x^5"

"..."

We note that


"x^n=F_nx+F_{n-1}"

Let the roots of our original quadratic be "u=\dfrac{1+\sqrt{5}}{2}," and "v=\dfrac{1-\sqrt{5}}{2}."

Since both "u" and "v" are roots of the quadratic, they must both satisfy 

"x^n=F_nx+F_{n-1}"

 So


"u^n=F_nu+F_{n-1}"

"v^n=F_nv+F_{n-1}"

Subtracting the second equation from the first equation yields


"u^n-v^n=F_n(u-v)"

Then


"(\dfrac{1+\sqrt{5}}{2})^{n+1}-(\dfrac{1-\sqrt{5}}{2})^{n+1}"

"=F_n(\dfrac{1+\sqrt{5}}{2}-\dfrac{1-\sqrt{5}}{2})"

"=\sqrt{5}F_n"

Then we have


"F_n=\dfrac{1}{\sqrt{5}}\bigg((\dfrac{1+\sqrt{5}}{2})^{n}-(\dfrac{1-\sqrt{5}}{2})^{n}\bigg)"


"F_{n-1}=\dfrac{1}{\sqrt{5}}\bigg((\dfrac{1+\sqrt{5}}{2})^{n-1}-(\dfrac{1-\sqrt{5}}{2})^{n-1}\bigg)"

"\dfrac{F_n}{F_{n-1}}=\dfrac{(\dfrac{1+\sqrt{5}}{2})^{n}-(\dfrac{1-\sqrt{5}}{2})^{n}}{(\dfrac{1+\sqrt{5}}{2})^{n-1}-(\dfrac{1-\sqrt{5}}{2})^{n-1}}"

"=\dfrac{\dfrac{1+\sqrt{5}}{2}-\dfrac{1-\sqrt{5}}{2}\cdot(\dfrac{1-\sqrt{5}}{1+\sqrt{5}})^{n-1}}{1-(\dfrac{1-\sqrt{5}}{1+\sqrt{5}})^{n-1}}"


"\lim\limits_{n\to \infin}\dfrac{F_n}{F_{n-1}}="

"=\lim\limits_{n\to \infin}\dfrac{\dfrac{1+\sqrt{5}}{2}-\dfrac{1-\sqrt{5}}{2}\cdot(\dfrac{1-\sqrt{5}}{1+\sqrt{5}})^{n-1}}{1-(\dfrac{1-\sqrt{5}}{1+\sqrt{5}})^{n-1}}"

"=\dfrac{\dfrac{1+\sqrt{5}}{2}-\dfrac{1-\sqrt{5}}{2}\cdot0}{1-0}"

"=\dfrac{1+\sqrt{5}}{2}"